Find the greatest angle through which a nonrelativistic alpha partiClec can be deflected by striking a stationary proton.[14°]


1
Expert's answer
2021-03-26T11:40:19-0400

Answer

the greatest angle through which a nonrelativistic alpha partiClec can be deflected by striking a stationary proton is given by

"tan\theta_1=\frac{sin\theta }{cos\theta +m_{\alpha}/m_p}\\=\frac{sin14° }{cos14° +4m_p/m_p}\\=\frac{0.24}{4.97}\\=0.048"

"\theta_1=2.75°"